# Weighted sums of Lucas numbers of the first and second kind

### v1

@@ 2 Preliminaries
Let $p$ and $q$ be nonzero (not necessarily coprime [^7] Ballot asserts that although Lucas himself and subsequently many other authors introduce the constraint that $\gcd(p,q)=1$, the mathematical literature contains many papers that prove results about the Lucas sequences with $\gcd(p,q)=1$, which hold just as well for Lucas sequences where $\gcd(p,q)\neq 1$ [1, p. 13].) integers. The Lucas sequences of the first and second kind, respectively, are denoted $\left(U_{n}(p,q)\right)_{n\geq 0}$ and $\left(V_{n}(p,q)\right)_{n\geq 0}$. Setting $U_{n}\coloneq U_{n}(p,q)$ and $V_{n}\coloneq V_{n}(p,q)$, the sequences $\left(U_{n}\right)_{n\geq 0}$ and $\left(V_{n}\right)_{n\geq 0}$ satisfy the recurrence relations $U_{n}=pU_{n-1}-qU_{n-2}$ and $V_{n}=pV_{n-1}-qV_{n-2}$, respectively, with initial conditions $U_{0}=0$, $U_{1}=1$, $V_{0}=2$, and $V_{1}=p$.
@@
$\left(V_{n}(2,-1)\right)_{n\geq 0}$ $=\left(Q_{n}\right)_{n\geq 0}$ companion Pell [^8] See Remark 2.4.
@@
$\left(V_{n}(6,1)\right)_{n\geq 0}$ $=\left(\mathfrak{C}_{n}\right)_{n\geq 0}$ double Lucas-balancing [^9] See Remark 2.5.
@@
$\left(V_{n}(3,2)\right)_{n\geq 0}$ $=\left(\mathcalb{F}_{n}\right)_{n\geq 0}$ Mersenne-Lucas. [^10] We denote a Mersenne-Lucas number with the symbol $\mathcalb{F}_{n}$ as it turns out that $\mathcalb{F}_{n}=2^{n}+1$, which are the so-called “Fermat numbers” whenever $n=2^{k}$ for some $k$.
@@ 3.1 Sum of consecutive terms when pq=1p-q=1
Recall that we are excluding the degenerate Lucas sequences as per Convention 2.3. But since we assume that pq=1p-q=1 in this subsection, this means we are excluding the cases when $(p,q)$ equals (0,1)(0,-1), $(1,0)$, or $(2,1)$. However, the following lemma holds even for the degenerate Lucas sequences.
@@ 4.1 Weighted sum of consecutive terms when pq=1p-q=1
where the second to last equality holds by Lemma 4.2, and the last equality holds since $(p,q)=(1,0)$ yields a degenerate sequence $\left(U_{n}(p,q)\right)_{n\geq 0}$ (recall Convention 2.3) and hence $q$ is nonzero. Thus the result follows. ∎

### v2

@@ 2 Preliminaries
Let $p$ and $q$ be nonzero (not necessarily coprime) integers. The Lucas sequences of the first and second kind, respectively, are denoted $\left(U_{n}(p,q)\right)_{n\geq 0}$ and $\left(V_{n}(p,q)\right)_{n\geq 0}$. Setting $U_{n}\coloneq U_{n}(p,q)$ and $V_{n}\coloneq V_{n}(p,q)$, the sequences $\left(U_{n}\right)_{n\geq 0}$ and $\left(V_{n}\right)_{n\geq 0}$ satisfy the recurrence relations $U_{n}=pU_{n-1}-qU_{n-2}$ and $V_{n}=pV_{n-1}-qV_{n-2}$, respectively, with initial conditions $U_{0}=0$, $U_{1}=1$, $V_{0}=2$, and $V_{1}=p$.
Remark 2.2.
Ballot and Williams assert that although Lucas and subsequently many other authors introduce the constraint that $\gcd(p,q)=1$, the mathematical literature contains many papers that prove results about the Lucas sequences with $\gcd(p,q)=1$, which hold just as well for Lucas sequences where $\gcd(p,q)\neq 1$ [1, p. 13].
@@
$\left(V_{n}(2,-1)\right)_{n\geq 0}$ $=\left(Q_{n}\right)_{n\geq 0}$ companion Pell (see Remark 2.5)
@@
$\left(V_{n}(6,1)\right)_{n\geq 0}$ $=\left(\mathfrak{C}_{n}\right)_{n\geq 0}$ double Lucas-balancing (see Remark 2.6)
@@
$\left(V_{n}(3,2)\right)_{n\geq 0}$ $=\left(\mathcalb{F}_{n}\right)_{n\geq 0}$ Mersenne-Lucas.
We denote a Mersenne-Lucas number with the symbol $\mathcalb{F}_{n}$ as it turns out that $\mathcalb{F}_{n}=2^{n}+1$, which are the so-called “Fermat numbers” whenever $n=2^{k}$ for some $k$.
@@ 3.1 Sum of consecutive terms when pq=1p-q=1
Recall that we are excluding the degenerate Lucas sequences as per Convention 2.4. But since we assume that pq=1p-q=1 in this subsection, this means we are excluding the cases when $(p,q)$ equals (0,1)(0,-1), $(1,0)$, or $(2,1)$. However, the following lemma holds even for the degenerate Lucas sequences.
@@ 4.1 Weighted sum of consecutive terms when pq=1p-q=1
where the second to last equality holds by Lemma 4.2, and the last equality holds since $(p,q)=(1,0)$ yields a degenerate sequence $\left(U_{n}(p,q)\right)_{n\geq 0}$ (recall Convention 2.4) and hence $q$ is nonzero. Thus the result follows. ∎